First Column:
C = 'completed'

Ranking:
A=4-5
B=3-4
C=2-3

LBC

S PI Object Filter Rank Req. time tot. exec. time UT date exec. time Notes
  SEMESTER 2010A
C Giallongo Q0933 G A 12000 13600 2010-01-17 13600 bad trasmission in part of the data
C Q0933 G A 12000 12000 2010-02-13 12000 redone because out of specs in Jan run
C Giallongo COSMOS G A 10000 2400 2010-01-17 2400 bad transmission in part of the data
C COSMOS G A 10000 1800 2010-02-13 1800 seeing ~ 1.0
C COSMOS G A 10000 10800 2010-02-14 10800 some images out of specs
  Giallongo Q1623 G A 7200 3600 2010-02-13 3600 probably out of specs
  Q1623 G A 7200 7200 2010-02-14 3600 probably out of specs
  Fiore COSMOS U A 43200 1800 2010-02-14 1800  
  COSMOS U A 43200 14400 2010-02-15 12600  
  COSMOS U A 43200 25200 2010-02-17 10800  
  COSMOS U A 43200 28800 2010-02-18 3600  
  COSMOS G A 3600 3600 2010-02-17 3600  
  COSMOS R A 3600 3600 2010-02-17 3600  
C Piotto M67 B B 7200 10800 2010-02-18 10800 some images out of specs: redone
C M67 I B 7200 10800 2010-02-18 10800 see B; binocular mode, i-SLOAN filter
C Rigliaco NGC2264-1 U B 1800 1800 2010-02-13 1800  
C NGC2264-1 B B 1800 1800 2010-02-13 1800  
C NGC2264-1 V B 1800 1800 2010-02-14 1800  
C NGC2264-1 R B 1800 1800 2010-02-14 1800  
C NGC2264-2 U B 1800 1800 2010-02-15 1800  
C NGC2264-1 B B 1800 1800 2010-02-15 1800  
C NGC2264-1 V B 1800 1800 2010-02-15 1800  
C NGC2264-1 R B 1800 1800 2010-02-15 1800  
C Palazzi GRB070714 U C 3600 2700 2010-01-17 2700 cloudy
C GRB070714 U C 3600 4500 2010-01-18 1800  
C GRB070714 B C 1800 1800 2010-01-17 1800  
C GRB070714 V C 1800 1800 2010-01-17 1800  
  Palazzi GRB070724 U C 3600        
C GRB070724 B C 1200 1200 2010-01-18 1800  
C GRB070724 V C 1200 1200 2010-01-18 1800  
  Palazzi GRB051222 U C 3600        
C GRB051222 B C 1800 1800 2010-01-18 1800  
C GRB051222 V C 1800 1800 2010-01-18 1800  
  Mereghetti LDRT Uspec A 14400 1800 2010-02-13 1800 FWHM ~1.2"
  LDRT Uspec A 14400 5400 2010-02-14 3600 slightly elongated?
  LDRT Uspec A 14400 9000 2010-02-15 3600 FWHM >~ 1.3"
  LDRT Uspec A 14400 10800 2010-02-17 1800  
  LDRT Uspec,i A 14400 12600 2010-02-18 1800 DX mirror hit limits but images are good
  Giallongo Q0933 r A 6000 5400 2010-04-11 5400 in most images the telescope is not guiding, probably out of specs
  Ferraro n5466 r A 9000 6400 2010-04-11 6400  
  n5466 B A 9000 6400 2010-04-11 6400  
  Giallongo Q1623 Uspec A 7200 ??? 9600 2010-06-13,15,16,18 7200 0.5h out of specs and 1OB repeated because of errors with Blue channel telemetry
  Q1623 z-Sloan A - 9600 2010-06-13,15,16,18 - not requested, observed since red channel available
  Ferraro n5466 r A 9000 4800 2010-06-13,16 2400 with this OB this program is completed
  n5466 B A 9000 4800 2010-06-13,16 2400 with this OB this program is completed
  Ferraro n5053 B A 9000 1560 2010-06-13 - OB out of specs ???
  n5053 r A 9000 1560 2010-06-13 - OB out of specs ???
C Rigliaco Serpens Uspec A ????? 7500 2010-06-13,16,17,18 6300 20min out of specs, high background
C Serpens R-bessel A ????? 3900 2010-06-13,17,18 3900  
C Serpens I-bessel A ????? 3600 2010-06-16,17,18 2400 20min out of specs, high background
C Serpens B-bessel A ????? 1500 2010-06-17,18 1500  
C Serpens V-bessel A ????? 1500 2010-06-17,18 1500  
  Piotto NGC 6791 V-bessel A 25200 ??? 6900 2010-06-15,16 6900 3000sec with red channel
  NGC 6791 I-bessel A 25200 ??? 6900 2010-06-15,16 6900  
  SEMESTER 2010B
  Clementini AndXXI-AndXIX B-bessel A 28800 5040 2010-10-11 5040 42min each out of specs
  AndXXI-AndXIX V-bessel A 28800 5040 2010-10-11 5040 42min each out of specs
  AndXXI-AndXIX B-bessel A 28800 5040 2010-10-11 5040 42min each out of specs
  AndXXI-AndXIX V-bessel A 28800 5040 2010-10-11 5040 42min each out of specs
  AndXXI-AndXIX B-bessel A 28800 11340 2010-11-30 6300  
  AndXXI-AndXIX V-bessel A 28800 11340 2010-11-30 6300  
  AndXXI-AndXIX B-bessel A 28800 17640 2010-12-01 6300  
  AndXXI-AndXIX V-bessel A 28800 17640 2010-12-01 6300  
  AndXXI-AndXIX B-bessel A 28800 23940 2010-12-02 6300  
  AndXXI-AndXIX V-bessel A 28800 23940 2010-12-02 6300  

LUCIFER

S PI Object Filter/grat Rank Req. time tot. exec. time UT date exec. time Notes
  SEMESTER 2010A
C Longhetti S2F1-145 150_K A 10800 9500 2010-01-16 9500 thin clouds
C S2F1-145 150_K A 10800 13100 2010-02-17 3600 clear and good seeing
C Mannucci Q1422 J A 3600 3600 2010-02-16 3600  
C Q1422 K A 3600 4500 2010-02-16 4500 serves also as MOS preimaging
  Q1422 210zJHK - K A 14400 10800 2010-04-09 10800 eventual problems with the calibration
  Q1422 210zJHK - K A 14400 14400 2010-04-10 3600 no Telluric done for the set
  Q1422 210zJHK - H A 14400 7200 2010-04-10 7200  
  Q1422 210zJHK - H A 14400 10200 2010-04-12 3000  
  Q1422 210zJHK - H A 14400 14700 2010-06-14 4500 arc-flat-tell
  Q1422 Ks A ???? 3600 2010-06-14 3600 to integrate with the preimaging for MOS
  Q1422 J A ???? 3600 2010-06-14 3600  
  HDF-N Ks A   900 2010-02-16 900 preimaging for MOS mask prep, 1 field
  HDF-N 210zJHK - K A 14400 7200 2010-04-12 7200  
  HDF-N 210zJHK - H A 14400 7200 2010-04-14 7200 the seeing very variable, maybe out of specs
  HDF-N 210zJHK - H A 14400 14400 2010-06-15 7200 arc-flat-tell
  HDF-N 210zJHK - K A 14400 10800 2010-06-16 3600 arc-flat-tell
  HDF-N 210zJHK - H A 14400 21600 2010-06-18 7200 arc-flat-tell
  HDF-N 210zJHK - K A 14400 12900 2010-06-18 2100 arc-flat-tell
C Turatto SN2009hv H C 7200 4800 2010-01-15 4800  
C SN2009kn 200HK C 7200 9000 2010-01-16 1500 bad transparency
  SN2009dc H C 7200 1800 2010-01-15 1800  
  Zamorani Cosmos Ks A   2700 2010-02-16 2700 preimaging for MOS mask prep, 3 fields
  Grazian GOODS-N Ks A   3600 2010-02-16 3600 preimaging for MOS mask prep, 4 fields
  GOODS-N1 200HK A 18000 7200 2010-04-09 7200  
  GOODS-N1 200HK A 18000 18000 2010-04-10 10800  
  GOODS-N 200HK A ???? 9000 2010-06-17 5400 no Telluric
  SEMESTER 2010B
  Longhetti S2F1-142 150Ks A 32400 31200 2010-10-09,10 32400 seeing below 1'' -Telluric+ARC+FLAT
  Turatto SN2009ig J, H, Ks B 640s+896s+896s 640 s + 896 s + 896 s 2010-10-07   seeing around 1"
  Turatto SN2010hq J, H, Ks B 20s+8s+8s 20s + 8s + 8s 2010-10-07   seeing below 1"
  SN2010hq 200HK A 800 800 2010-10-07 800 seeing below 1" - Telluric + ARC +FLAT
  Turatto SN2010gi J, H, Ks B 40+80+80 40+80+80s 2010-10-07 800 seeing around 1"
  SN2010gi 200HK A 3680 3680 2010-10-08 3680 seeing below 1" - Telluric+ARC+FLAT
  SN2010gi J, H, Ks A   80 s + 160 s +240 s low airmass
  Turatto SN2009jf J, H, Ks B 80+160+240 80+160+240 2010-10-08 80+160+240 seeing around 1"
  Turatto SN2009kr J, H, Ks B   640 s + 896 s + 896 s 2010-10-10    
  SN2009kr J, H, Ks A   640 s + 896 s + 896 s 2010-11-30    
  Turatto SN2010al J, H, Ks B   640 s + 896 s + 896 s 2010-10-10    
  Turatto SN2010jl J, H, Ks A   12 s + 12 s + 24 s 2010-11-30   guiding probe covers 1.4 of field during 3rd offset
  SN2010jl 200HK A   2000sec 2010-11-30   Tell+ARC+FLAT
  SN2010jl J, H, Ks A   12 s + 12 s + 24 s 2010-12-01   repeated with different guide star
  Turatto SN2009kn J, H, Ks A   640 s + 896 s + 896 s 2010-12-01    
  Cresci DSF2237A Ks A 14400 3000 2010-10-08 3000 seeing below 1", maybe 5min with clouds
  DSF2237a Ks A 14400 4380 2010-10-09 7380  
  Cresci DSF2237b Ks A 14400 4500 2010-10-09 4500  
  Cresci HDF 210zJHK+H A 6hr 1hr 2010-11-30   ARC+FLAT - NO Tell - use the one in Dec02
  HDF 210zJHK+H A 6hr 1hr 2010-12-01   ARC+FLAT - Tell
  HDF 210zJHK+K A 2hr 1hr 2010-12-01   ARC+FLAT - Tell
  HDF 210zJHK+H A 6hr 1hr25min 2010-12-02   ARC+FLAT - Tell
  HDF 210zJHK+K A 2hr 1hr 2010-12-02   ARC+FLAT - Tell
  Pentericci Cluster1 200H+K A 18000 3600 2010-10-09 3600 data may be compromised
  cluster1HK 200HK A 18000 3600 2010-12-02 7200 from airmass 1.5 to 3
  cluster1HK 200HK A 18000 15600 2010-12-02 22800 some clouds, must be verified

-- FilippoMannucci - 18 Jan 2010 -- VincenzoTesta - 03 Dec 2010
Topic revision: r14 - 04 Dec 2010, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback